4. What initial volume of solution is needed to dilute a 1 g/ml solution to a 1 μg/ml solution? The desired final volume solution is 20 ml.


5. The initial 50 mL solution is 40 M concentrated. Now I want a 2 M concentrated solution instead. How can I accomplish this?
Answer #1

4. C1V1 = C2V2

1 × V1 = 10^-6 × 20

V1 = 20 × 10^-6 = 20 ul

20 ml = 20 × 1000 = 20,000 ul

Volume of 1 g/ml solution required = 20 ul

Volume of diluent = 20,000 - 20 = 19980 ul = 19.980 ml

5. C1V1 = C2V2

40 × 50 = 2 × V2

V2 = 20 × 50 = 1000 ml

Final volume should be 1000 ml. So, to 50 ml of 40 M solution, add diluent and make final volume upto 1000 ml (1 L).

Please rate high.